f(x)=3x^2-2x
f(0)=3(0)^2-2(0)=0
likewise for f-(-1), let x=-1 f(-1)=5 i don't undestand how they came up with 5
finally (2)=2(2)^2-2(2)=8

Answer by richard1234(7193)

f(-1) = 5 because f(-1) = 3(-1)^2 - 2(-1) = 3 + 2 = 5.